# GeMID: Generalizable Models for IoT Device Identification

### v1

@@ IV-B Feature Interactions
We then considered feature interactions, using a wrapper method, specifically a Genetic Algorithm (GA), to find the optimal combination of individually predictive features which we validated across multiple datasets. The global search characteristic of a GA tends to find good solutions, and avoiding exhaustive search enables us to obtain potential feature sets within an acceptable timeframe. However, in general there is no guarantee of a GA finding an optimal subset, meaning it may include uninformative features in the feature set. Additionally, GA-selected features may develop a bias towards the data used for selection, potentially reducing their performance on other datasets.
@@
To address this, all 46 features that passed individual voting were combined into a single feature set for selection using a genetic algorithm (GA). A decision tree (DT) model was employed for evaluation, with the fitness of each generation assessed using the F1 score across 8 additional DD datasets (see Fig. 4). This external validation provides feedback to the GA, ensuring that feature selection is based on success across different datasets and not dependent on only one. For each of the eight DD cases, the GA produced slightly different feature sets, so we then identified the intersections between them. Fig. 7 depicts this, showing that some features (e.g., tcp.window_size, tcp.ack) were chosen multiple times, some (e.g., tcp.seq, tcp.flags.ack) only once.
IV-C Deriving a Final Feature Set
@@
To identify the most effective combination of features from the GA results, we conducted two post hoc analyses. First, the feature set derived from each DD case (each GA run) is reevaluated on each other DD case. F1 scores of the resulting DT models are presented in the left part of Fig. 8. These show how well the feature sets from each GA run generalize.
@@
The second analysis uses a voting mechanism based on intersections of the eight GA feature sets (Fig. 7), with a feature receiving votes proportional to how often it appears. Features are then grouped based on thresholds, i.e., whether they appear in 2 or more feature sets (Vote+2), 3 or more (Vote+3) etc., and DT models are trained using these feature groups. The results are shown on the right side of Fig 8.
@@
From the mean F1 scores shown in the final row of Fig. 8, it can be seen that feature sets based on grouping generally lead to better performance, at least for voting thresholds up to 4. For larger thresholds, the feature sets become very small, explaining the lower scores. Generally, the feature sets from single runs lead to models with lower F1 scores, suggesting that voting adds more robustness. Consequently, we use the Vote+3 feature set — which has the highest F1 score — in the next section. The selected features are highlighted in Fig. 7.
@@ V-A Model Selection
We used random search for hyperparameter optimization and applied each model to the DD datasets — see Supplementary Material (SM) Section 4.4 for details. The F1 scores and average inference times are shown in Fig. 9. From this, it is clear that the most successful models are RF and XGB, with F1 scores of 0.799 and 0.780, respectively. While their predictive performance is similar, RF runs about 6 times faster than XGB, so we chose this as our preferred model. Notably, DT has the fastest inference time, but its predictive performance lags behind RF by about 10 percentage points.
@@
V-B Evaluating Model Generalizability
In this section, we measure the generalizability of our method (GeMID) and other methods using the MonIoTr dataset, using the US, US-VPN, UK, and UK-VPN partitions of the MonIoTr dataset within three different evaluation contexts. In CV, we carry out cross-validation within a single partition. In SS, we train models using non-VPN data and test them using VPN data, and vice versa. In DD, we use data collected from one geographical site to train models, and data collected from the other geographical site to test them.
As comparison baselines, we use flow-based features generated by CICFlowmeter [30] and window-based features generated by Kitsune [31]. These are well-known tools that were used in previous DI studies. To ensure a fair comparison, feature and model selection were repeated independently (see SM Section I and II for results). We also compare against IoTDevID [13], another packet header-based DI method that aimed to build generalizable DI models and demonstrated improved performance over alternative methods.
TABLE I: Comparison of DI methods on MonIoTr dataset. F1 scores.
@@ V-B Evaluating Model Generalizability
DD UK|US 0.760 0.475 0.703 0.464
UK|USVPN 0.769 0.490 0.666 0.435
UKVPN|US 0.748 0.516 0.654 0.392
UKVPN|USVPN 0.746 0.454 0.709 0.444
US|UK 0.778 0.454 0.752 0.541
US|UKVPN 0.780 0.517 0.694 0.455
USVPN|UK 0.811 0.515 0.710 0.439
USVPN|UKVPN 0.816 0.464 0.731 0.521
@@
The results are shown in Table I. All methods achieved high scores for CV, close to or above 0.90. However, a decrease is observed in all methods for SS and DD, with this decrease being particularly dramatic for the statistics-based methods (CICFlowMeter and Kitsune). For DD, GeMID maintains 81% of its CV F1 score and IoTDevID 77%, compared to only 55% for CIC and 50% for Kitsune. This significant drop supports our claim that flow or window-based statistical methods, which focus on relationships between devices within the network rather than individual device activities, are not suitable for building generalizable DI models due to their poor transferability across heterogeneous network environments.
The two methods — GeMID and IoTDevID — based on packet headers both perform significantly better, with GeMID outperforming IoTDevID within all evaluation contexts. Table II shows the per-device results for GeMID within the DD evaluation context. It can be seen that most of the devices are identified with a high level of accuracy. A notable exception is the echo family of devices, which, presumably due to underlying similarities, are easily misclassified with each other (see SM Figs. 12 and 13 for confusion matrices). There is also a degree of misclassification between devices from different manufacturers that perform the same function, e.g., between roku-tv and samsungtv-wired, and between yi-camera and wansview-cam-wired. A third category of poor performance is due to data paucity, e.g., for sousvide, there are only 70 US samples, compared to 15,000 for UK. The challenges in classifying these devices are not specific to GeMID but are also observed in other methods (IoTDevID, CICFlowMeter, and Kitsune). These methods also encounter similar difficulties due to the inherent complexity of distinguishing between devices with similar characteristics.
TABLE II: F1 scores per device in all 8 DD cases for MonIoTr data, with green for higher and red for lower success.
Devices Train→ UK UK UKVPN UKVPN US US USVPN USVPN
Test→ US USVPN US USVPN UK UKVPN UK UKVPN
@@ V-C Feature Usage
Notably, IoTDevID’s feature set is rich in features related to BOOTP, DNS, and EAPOL protocols, whereas GeMID lacks these. This may be due to a dataset-specific bias, owing to the reliance on a single dataset — Aalto — for IoTDevID’s feature selection process. This is supported by the observation that while GeMID outperforms IoTDevID in experiments using the UNSW and MonIoTr datasets, IoTDevID shows markedly better performance when applied to the Aalto dataset [13] (see SM Table VIII).
@@
Another reason for the difference in performance may be the larger initial pool of extracted features in GeMID (332 features) compared to IoTDevID (112 features). Although both studies generally use the same protocols, GeMID captures more sub-features within each protocol, providing finer-grained information in the feature set. In terms of selected features, GeMID offers more comprehensive protocol coverage, starting with a richer set of HTTP features and extending to broader coverage of TCP and UDP protocols. This includes features such as sequence numbers, timing metrics, and window size information, which enhance its ability to capture more detailed network behavior (see Fig. 7). Unlike in the IoTDevID study, we retained features containing session-based identifiers. Whilst these can lead to over-optimistic metric scores [32], this is only the case when sessions are split between test and train sets, e.g., when using CV. When this is controlled for using stricter evaluation contexts (as in this study), the inclusion of these features can enhance model performance by capturing meaningful relationships between packets.
@@
In practice, the accuracy of packet-level classification methods can be improved by aggregating predictions from individual packets. The IoTDevID study introduced an aggregation algorithm that enhances performance by combining multiple packet-level predictions. This approach is particularly effective for packet-level methods such as GeMID, IoTDevID, and Kitsune (despite Kitsune focusing on the relationship between the window system and packets, as it still relies on packet-based labels). In contrast, CICFlowMeter operates at the flow level, where such aggregation is not applicable. Our results in SM Tables VI and VII show the impact of this approach, with GeMID achieving a mean F1 score of 0.892 in the DD evaluation context.
VI Limitations
@@ VII Conclusions
Comparative analysis of DI methods showed that packet header-based features performed better than network statistics and window methods. Packet features offer a more consistent and reliable basis for model training because they are less influenced by network-specific variables. This finding highlights the importance of selecting features that inherently capture device-specific attributes.

### v2

@@ 4.2 Feature Interactions
However, in general there is no guarantee of a GA finding an optimal subset, meaning it may include uninformative features in the feature set. Additionally, GA-selected features may develop a bias towards the data used for selection, potentially reducing their performance on other datasets.
@@
To address this, all 46 features that passed individual voting were combined into a single feature set for selection using a genetic algorithm (GA). A decision tree (DT) model was employed for evaluation, with the fitness of each generation assessed using the F1 score across 8 additional DD datasets (see Figure 4). This external validation provides feedback to the GA, ensuring that feature selection is based on success across different datasets and not dependent on only one. For each of the eight DD cases, the GA produced slightly different feature sets, so we then identified the intersections between them. Figure 7 depicts this, showing that some features (e.g., tcp.window_size, tcp.ack) were chosen multiple times, some (e.g., tcp.seq, tcp.flags.ack) only once.
4.3 Deriving a Final Feature Set
@@
To identify the most effective combination of features from the GA results, we conducted two post hoc analyses. First, the feature set derived from each DD case (each GA run) is reevaluated on each other DD case. F1 scores of the resulting DT models are presented in the left part of Figure 8. These show how well the feature sets from each GA run generalize.
@@
The second analysis uses a voting mechanism based on intersections of the eight GA feature sets (Figure 7), with a feature receiving votes proportional to how often it appears. Features are then grouped based on thresholds, i.e., whether they appear in 2 or more feature sets (Vote+2), 3 or more (Vote+3) etc., and DT models are trained using these feature groups. The results are shown on the right side of Figure 8.
@@
From the mean F1 scores shown in the final row of Figure 8, it can be seen that feature sets based on grouping generally lead to better performance, at least for voting thresholds up to 4. For larger thresholds, the feature sets become very small, explaining the lower scores. Generally, the feature sets from single runs lead to models with lower F1 scores, suggesting that voting adds more robustness. Consequently, we use the Vote+3 feature set—which has the highest F1 score—in the next section. The selected features are highlighted in Figure 7.
@@ 5.1 Model Selection
We used random search for hyperparameter optimization and applied each model to the DD datasets—see Appendix Section E.4 for details. The F1 scores and average inference times are shown in Figure 9. From this, it is clear that the most successful models are RF and XGB, with F1 scores of 0.799 and 0.780, respectively. While their predictive performance is similar, RF runs about 6 times faster than XGB, so we chose this as our preferred model. Notably, DT has the fastest inference time, but its predictive performance lags behind RF by about 10 percentage points.
This performance difference is characteristic of the trade-off between single learners and ensemble models. A single DT is computationally simple and therefore very fast, but it can be prone to overfitting the training data. RF mitigates this by building a multitude of decision trees on different subsets of the data and features, and then averaging their predictions. This ensemble approach reduces the model’s overall variance and improves its ability to generalize to unseen data, resulting in a higher F1 score at the cost of increased computational overhead for inference.
@@
5.2 Evaluating Model Generalizability
In this section, we measure the generalizability of our method (GeMID) and other methods using the MonIoTr dataset, using the USA, USA-VPN, UK, and UK-VPN partitions of the MonIoTr dataset within three different evaluation contexts. In CV, we carry out cross-validation within a single partition. In SS, we train models using non-VPN data and test them using VPN data, and vice versa. In DD, we use data collected from one geographical site to train models, and data collected from the other geographical site to test them.
To test our hypothesis about the fragility of statistical features, we needed to compare our packet-based approach against strong, representative baselines for flow-based and window-based methods. For this purpose, we selected CICFlowmeter [38] and Kitsune [39] not as published methods to compare against, but as well-established tools for extracting comprehensive sets of flow and window-based features, respectively. These tools are powerful and generate a wide array of statistical features, making them ideal for creating robust baseline models that effectively represent the statistical methods prevalent in the DI literature. We then applied our own rigorous feature selection and evaluation pipeline to these feature sets to ensure a fair and direct comparison of the feature types themselves.
To ensure a fair comparison, feature and model selection were repeated independently (see Appendix Sections B and C for results). We also compare against IoTDevID [13], another packet header-based DI method that aimed to build generalizable DI models and demonstrated improved performance over alternative methods.
Table 1: Comparison of DI methods on MonIoTr dataset, in F1 scores. The final section of the table quantifies the average generalization gap ($\Delta F1$) between the different evaluation contexts (Cross-Validation, Session-vs-Session, and Dataset-vs-Dataset), where a smaller value indicates better generalizability.
@@ 5.2 Evaluating Model Generalizability
DD UK|USA 0.760 0.475 0.703 0.464
UK|USAVPN 0.769 0.490 0.666 0.435
UKVPN|USA 0.748 0.516 0.654 0.392
UKVPN|USAVPN 0.746 0.454 0.709 0.444
USA|UK 0.778 0.454 0.752 0.541
USA|UKVPN 0.780 0.517 0.694 0.455
USAVPN|UK 0.811 0.515 0.710 0.439
USAVPN|UKVPN 0.816 0.464 0.731 0.521
@@
$\Delta$F1 CV-SS 0.079 0.274 0.088 0.213
CV-DD 0.185 0.393 0.208 0.464
SS-DD 0.106 0.119 0.12 0.251
The results are shown in Table 1. All methods achieved high scores for CV, close to or above 0.90. However, a decrease is observed in all methods for SS and DD, with this decrease being particularly dramatic for the statistics-based methods (CICFlowMeter and Kitsune). For DD, GeMID maintains 81% of its CV F1 score and IoTDevID 77%, compared to only 55% for CIC and 50% for Kitsune. This significant drop supports our claim that flow or window-based statistical methods, which focus on relationships between devices within the network rather than individual device activities, are not suitable for building generalizable DI models due to their poor transferability across heterogeneous network environments.
The performance drop between evaluation contexts can be more formally analyzed by examining the generalization gap ($\Delta F1$), summarized in the final section of Table 1. This metric quantifies the loss in F1 score as the testing conditions become more challenging. The most critical gap is between the lenient cross-validation and the strict dataset-versus-dataset context (CV-DD). Here, the packet-header-based methods, GeMID and IoTDevID, show superior resilience with gaps of 0.185 and 0.208, respectively. In contrast, the statistics-based methods fail to generalize effectively, with CICFlowMeter exhibiting a gap of 0.393 and Kitsune a gap of 0.464—more than double that of GeMID. This quantitatively demonstrates that statistical features are not robust to changes in the network environment. A more detailed, per-case analysis of this generalization gap is available in Appendix Table 16.
The two methods—GeMID and IoTDevID—based on packet headers both perform significantly better, with GeMID outperforming IoTDevID within all evaluation contexts. Table 2 shows the per-device results for GeMID within the DD evaluation context. It can be seen that most of the devices are identified with a high level of accuracy. A notable exception is the echo family of devices, which, presumably due to underlying similarities, are easily misclassified with each other (see Appendix Figures 21 and 22 for confusion matrices). There is also a degree of misclassification between devices from different manufacturers that perform the same function, e.g., between roku-tv and samsungtv-wired, and between yi-camera and wansview-cam-wired. A third category of poor performance is due to data paucity, e.g., for sousvide, there are only 70 USA samples, compared to 15,000 for UK. The challenges in classifying these devices are not specific to GeMID but are also observed in other methods (IoTDevID, CICFlowMeter, and Kitsune). These methods also encounter similar difficulties due to the inherent complexity of distinguishing between devices with similar characteristics.
Table 2: F1 scores per device in all 8 DD cases for MonIoTr data, with green for higher and red for lower success.
Devices Train→ UK UK UKVPN UKVPN USA USA USAVPN USAVPN
Test→ USA USAVPN USA USAVPN UK UKVPN UK UKVPN
@@ 5.3 Feature Usage
Notably, IoTDevID’s feature set is rich in features related to BOOTP, DNS, and EAPOL protocols, whereas GeMID lacks these. This may be due to a dataset-specific bias, owing to the reliance on a single dataset—Aalto—for IoTDevID’s feature selection process. This is supported by the observation that while GeMID outperforms IoTDevID in experiments using the UNSW and MonIoTr datasets, IoTDevID shows markedly better performance when applied to the Aalto dataset [13] (see Appendix Table 8).
@@
Another reason for the difference in performance may be the larger initial pool of extracted features in GeMID (332 features) compared to IoTDevID (112 features). Although both studies generally use the same protocols, GeMID captures more sub-features within each protocol, providing finer-grained information in the feature set. In terms of selected features, GeMID offers more comprehensive protocol coverage, starting with a richer set of HTTP features and extending to broader coverage of TCP and UDP protocols. This includes features such as sequence numbers, timing metrics, and window size information, which enhance its ability to capture more detailed network behavior (see Figure 7). Unlike in the IoTDevID study, we retained features containing session-based identifiers (such as TCP acknowledgment, TCP sequence, and source/destination port numbers). Whilst these can lead to over-optimistic metric scores [40], this is only the case when sessions are split between test and train sets, e.g., when using CV. When this is controlled for using stricter evaluation contexts (as in this study), the inclusion of these features can enhance model performance by capturing meaningful relationships between packets.
@@
In practice, the accuracy of packet-level classification methods can be improved by aggregating predictions from individual packets. The IoTDevID study introduced an aggregation algorithm that enhances performance by combining multiple packet-level predictions. This approach is particularly effective for packet-level methods such as GeMID, IoTDevID, and Kitsune (despite Kitsune focusing on the relationship between the window system and packets, as it still relies on packet-based labels). In contrast, CICFlowMeter operates at the flow level, where such aggregation is not applicable. Our results in Appendix Tables 9 and 10 show the impact of this approach, with GeMID achieving a mean F1 score of 0.892 in the DD evaluation context.
6 Limitations and Practical Considerations
6.1 Practical Considerations and Deployment Scenarios
While the core of this research is methodological, it is important to consider the practical path to deployment for a DI system built on our framework.
Deployment Environment: We envision the GeMID classifier operating at the edge of the local network, for instance, on a home router or an enterprise gateway. This location is ideal as it has visibility into all traffic from local IoT devices and allows for device-specific policies (e.g., isolation, access control) to be enforced locally. Deployment at the ISP level, while possible, would present significant privacy and scalability challenges.
Model Scalability and Management: The multi-class models developed in this paper were aimed at assessing the effectiveness of our feature set across a predefined list of devices. In a real-world, dynamic environment, a more practical approach would be a combination of binary, one-vs-all models. When a new device is added to the network, the system could either attempt to classify it against its known models or flag it as unknown. If a new model for this device type becomes available, it can be added to the system without retraining the entire multi-class classifier. This binary model approach inherently solves the issue of adding or removing devices from a household.
Model Training and Distribution: The responsibility for training would likely fall to device manufacturers, who are best positioned to generate traffic data for their products under controlled conditions. These pre-trained, signed models could be made available through a central repository. A network controller, perhaps using Software-Defined Networking (SDN) principles, could then automatically detect a new, unclassified device, query the repository for a corresponding model, and integrate it into the local DI system.
Model Size and Efficiency: A key concern for deployment on resource-constrained hardware like a home router is model efficiency. Our framework directly addresses this. The final GeMID model utilizes a Random Forest classifier with a small, optimized feature set. As shown in our evaluation (Figure 9), this model has a very fast average inference time, making it highly efficient and suitable for online, real-time deployment without significant computational overhead.
6.2 Study Limitations and Future Work
@@
A second limitation relates to GeMID’s dependence on packet header features. Our method relies on features extracted from unencrypted headers such as those from IP, TCP, and UDP. This approach remains effective for a majority of today’s IoT traffic where the payload is encrypted by a protocol like TLS but the transport headers remain visible. However, the increasing adoption of newer protocols, most notably QUIC, which encrypts most of the transport layer header information, presents a future challenge. The effectiveness of GeMID’s current feature set would likely be diminished on networks with heavy QUIC adoption, and future work should investigate features that are robust to transport-layer encryption.
Third, it is important to discuss the computational overhead of our framework. The GA used for feature selection is a computationally intensive process. However, this is a one-time, offline cost incurred during the model design and training phase. It is not part of the operational, real-time device identification system. The final GeMID model, which uses a small, optimized feature set with a Random Forest classifier, is highly efficient and suitable for online deployment, as evidenced by the fast inference times shown in Figure 9. Therefore, while the development phase requires significant resources, the deployment phase is lightweight.
Finally, the scalability of GeMID to environments with a significantly larger number of distinct device types has not yet been validated. Our experiments were conducted on datasets with several dozen unique devices. Scaling to networks with hundreds or thousands of device classes would make the multi-class classification task inherently more complex. This could potentially reduce classification accuracy, especially among devices with very similar functionalities and traffic patterns. Therefore, while our framework is designed to produce generalizable models, its performance limits at a very large scale remain an open question for future research.
@@ 7 Conclusions
Comparative analysis of DI methods showed that packet header-based features performed better than network statistics and window methods for building generalizable DI models. Packet features offer a more consistent and reliable basis for model training because they are less influenced by network-specific variables. This claim is most strongly supported by our generalization gap analysis (Table 1), where the performance of statistical methods like CICFlowMeter and Kitsune degraded by over 100% more than our packet-based GeMID method when moving from single-environment validation to cross-environment testing (CV-DD $\Delta$F1 of 0.393 and 0.464 vs. 0.185). This demonstrates that packet features offer a more consistent and reliable basis for model training because they capture device-intrinsic attributes rather than network-specific variables. This finding highlights the importance of selecting features that inherently capture device-specific attributes.